### v1

# thanks: $0$-0002-5209-872Xthanks: $0$-0002-1235-0505

Abstract
The next generation of neutrino oscillation experiments, JUNO, DUNE, and HK, are under construction now and will collect data over the next decade and beyond. As there are no approved plans to follow up this program with more advanced neutrino oscillation experiments, we consider here one option that had gained considerable interest more than a decade ago: a neutrino factory. Such an experiment uses stored muons in a racetrack configuration with extremely well characterized decays reducing systematic uncertainties and providing for more oscillation channels. Such a machine could also be one step towards a high energy muon collider program. We consider a long-baseline configuration to SURF using the DUNE far detectors or modifications thereof, and compare the expected sensitivities to the three-flavor oscillation parameters to the anticipated results from DUNE and HK. We find that a neutrino factory can improve our understanding of CP violation and also aid in disentangling the complicated flavor puzzle.
I Introduction
The next generation of neutrino oscillation experiments, including JUNO Abusleme et al. (2022), DUNE Abi et al. (2020), and HK Abe et al. (2015) are expected to make significant headway in addressing three of the remaining known unknowns in particle physics. Notably, they will have excellent $\gg 5\sigma$ sensitivity to the mass ordering (the sign of $\Delta m^{2}_{31}$), they will have good $\theta_{23}$ octant determination capabilities provided that it isn’t too close to maximal, and both DUNE and HK will each individually provide $>5\sigma$ sensitivity to CP violation for significant regions of parameter space. In addition, they will provide some resolution on measuring $\delta$ (we use the standard parameterization of the PMNS Pontecorvo (1957); Maki et al. (1962) matrix Denton and Pestes (2021)). Atmospheric neutrino experiments such as HK, the IceCube-upgrade Ishihara (2021), and KM3NeT-ORCA Adrian-Martinez et al. (2016) will also provide complementary information. For a recent overview of this picture, see Denton et al. (2022).
While the outlook is hopefully rosy, it is important to carefully examine what, if anything, should come after these experiments in the neutrino oscillation program. Several experiments have been proposed with varying levels of detail. One is using a water-based liquid scintillator detector THEIA which combines the advantages of water and liquid scintillator, likely in place of one of the four LArTPC modules for DUNE Askins et al. (2020, 2022). Another is an additional HK-like tank in Korea experiencing a different portion of the same beam that HK experiences to target the second oscillation maximum in appearance Abe et al. (2018). There is also discussion about building a very large (much larger than HK) water Cherenkov detector in Sweden combined with additional beam upgrades of the ESS to create the ESSnuSB also to target the second oscillation maximum in Alekou et al. (2021, 2022); Abele et al. (2023).
In this article we discuss a different option, a neutrino factory. Neutrino factories were seriously considered as an option to probe CP violation in the first decade of this millennium largely due to the expectation of small $\theta_{13}<1^{\circ}$ which severely limits the possible size of the CP signal De Rujula et al. (1999). In general, a neutrino factory is a muon storage ring in a racetrack configuration with long straight sections pointing towards the far detector and provides powerful beams of both muon and electron neutrinos of opposite charges Geer (1998). The initial spectrum of this decay is extremely well characterized unlike the initial spectrum from traditional fixed target neutrino oscillation experiments which primarily produce neutrinos from light meson decays with a range of energies. See Bogacz et al. (2022) for a recent review of the history of neutrino factories as well as a discussion of the non-oscillation physics program at such a machine.
A neutrino factory may also make sense as a first step on the path to a multi-TeV high energy muon collider to probe the electroweak sector, see e.g. Accettura et al. (2023). This first step could be chronologically as a part of a technology demonstrator, a part of the accelerator chain, or both. In addition, the recent 2023 P5 report suggested consideration of neutrino physics to be done with a muon accelerator p5_ . A neutrino factory may also be a necessary successor to the upcoming experiments if there is a disagreement in the determination of the oscillation parameters among the different experiments; a situation that the neutrino field is well accustomed to.
In this paper we will discuss the oscillation physics at a neutrino factory considering the current knowledge of the oscillation parameters and address how a modern look at the oscillation physics case for a neutrino factory compares to previous studies. We then derive numerical sensitivities for the precision on $\delta$ for a potential setup of a neutrino factory for two different baselines and energies, before we conclude.
II Neutrino Factory Setup
At a neutrino factory the neutrinos originate from the decays of positively or negatively charged muons, leading to a beam of (anti-)electron neutrinos and (anti-)muon neutrinos. There are a few main differences for the neutrino beam of a neutrino factory compared to conventional neutrino beams from fixed target experiment:
  • The achievable maximal neutrino energy is higher at a neutrino factory as the produced muon neutrinos have energies up to the muon energy. Indeed the muon neutrino energy spectrum rises with energy instead of falling like at a fixed target neutrino oscillation experiment, see fig. 1.
  • The composition and the expected energy of the neutrino beam is well known as radiative corrections to muon decay are small Greub et al. (1994); Laing and Soler (2008); Tomalak (2022). Furthermore, no $\nu_{\tau}$ get produced in the source, allowing for clean searches for $\nu_{\tau}$ appearance channels.
  • The neutrino beam at a neutrino factory is composed of equally many neutrinos as anti-neutrinos, albeit of different flavors. In particular, the flux consists of 50% $\nu_{e}$ which allows for $\nu_{e}\to\nu_{\tau}$ searches, unlike at fixed target oscillation experiments where the $\nu_{e}$ flux is designed to be small.
  • The neutrino energy is tunable and flexible, which allows for physics studies at different, well-known energies.
The first studies of a neutrino factory studies were focused on measuring a small $\theta_{13}$, $\delta$, and the mass ordering and identified problems with degeneracies in the oscillation analysis Barger et al. (2002); Freund et al. (2001); Huber et al. (2006). For small values of $\theta_{13}$, charge ID (CID) of muons or electrons was crucial to suppress the beam background for the appearance channels and to measure $\delta$. For this reason, many studies assumed a 100 kT magnetized iron detector Choubey et al. (2011).
With the current and envisioned knowledge of all oscillation parameters and the present experimental landscape, the task of a future neutrino factory is no longer to discover the oscillation parameters but more likely to make precise measurements and/or potentially resolve any discrepancies identified in previous measurements. Therefore, with a different physics goal and a more modern experimental landscape, the experimental requirements on a neutrino factory also necessarily evolve.
Figure 1: The initial neutrino flux composition at DUNE and a neutrino factory from protons on a fixed target (left axis) or muon decays at the energy optimized for the FNAL-SURF baseline: $E_{\mu}=5$ GeV (right axis).
We consider two possible neutrino factory scenarios: the neutrino source is Fermilab and, to make use of existing facilities, the far detector to be located at SURF, resulting in a baseline of 1284.9 km. Alternatively, we also consider the neutrino source at Brookhaven National Laboratory, to make use of the AGS/RHIC/EIC accelerator complex, which leads to baseline of 2542.3 km to SURF. We assume as matter density for both configurations 2.848 g/$\text{cm}^{3}$. We assume that the far detector is based on a LArTPC configuration with a total fiducial target mass of 40 kT consistent with the envisioned DUNE detector. Based on previous exploratory studies of neutrino factories we assume $10^{21}$ muon (or anti-muon) decays per year Choubey et al. (2011).
To simulate a neutrino factory we make use of DUNE’s far detector design Abi et al. (2021) in the GLoBES Huber et al. (2005) framework. We also consider oscillated tau neutrinos events with a tau neutrino CC cross section from Yaeggy (2023). We include tau events as background only Donini et al. (2011) and don’t consider tau signal events for this three-flavor oscillation study due to the large uncertainties on the signal efficiency at LAr detectors; including this channel would improve the sensitivity to measure the oscillation parameters, although likely at a subleading level due to the lower statistics because the $\nu_{\tau}$ cross section is suppressed at these energies. As the initial neutrino flux is well known we assume a normalization uncertainty of 2% on the electron signal and background, 5% on the muon signal and background, 10% on the NC background, 20% on the tau background.
III Sensitivities
For our simulations we consider priors on the oscillation parameters other than $\delta$ from Gonzalez-Garcia et al. (2021) and we fix the MO to NO, as the MO as well as all oscillation parameters will be determined via multiple independent methods with future experiments and our results are largely independent of the true MO. Therefore we do not study the sensitivity of a neutrino factory on these parameters but instead study the possible improvements a neutrino factory can bring on the precision on $\delta$. This provides an estimate of the ability of a neutrino factory to differentiate among differing measurements from upcoming experiments.
We studied the sensitivity to the oscillation parameters as a function of muon energy as well as the ratio of muon to anti-muon running time. For the FNAL-SURF (BNL-SURF) baseline we find that a muon energy of $E_{\mu}=5$ (8) GeV with equal muon and anti-muon running time allows for the most precision on the determination of $\delta$ for most cases. The higher energy is necessary for the longer baseline to get the appearance maximum to occur within the muon decay spectrum at that baseline. Depending on the specific details of the experimental configuration as well as progress in measuring the oscillation parameters at NOvA and T2K, SK and IceCube, as well as DUNE and HK, may lead to slight modifications on these optimal energies.
In fig. 2 we show the expected $1\sigma$ precision on $\delta$ as a function of CID and exposure for the FNAL-SURF setup for $\delta_{\text{true}}=0,-90^{\circ}$. We include the anticipated information from 10 years each of DUNE and HK. For DUNE we assume 480 kT-MW-year which corresponds to 5 years of each neutrino running and anti-neutrino with 1.2 MW proton beam and with a total fiducial volume of 40 kT of LAr. For HK we assume 190 kT water detector, 1.3 MW beam running for 10 years with $\nu:\bar{\nu}=1:3$ from Scott (2021). We simulate both experiments within the GLoBES framework.
Since the beam at a neutrino factory consists of equal amounts of neutrinos and anti-neutrinos, there is an inherent large contribution to the final states coming from the other charge, i.e. the processes $\nu_{e}\to\nu_{\mu}$ and $\bar{\nu}_{\mu}\to\bar{\nu}_{\mu}$ lead to the same final state although with a different energy spectrum due to different initial spectra and different oscillation probabilities. Differentiating the events from the disappearance channel in the analysis of the disappearance channel has been identified as an important step towards a successful neutrino factory early on De Rujula et al. (1999). Here we study the impact of CID efficiency $\epsilon_{CID}$ in terms of either electron or muon charge identification. We define the CID parameter $\epsilon_{CID}$ as
$N_{\nu_{\text{f,obs}}}$ $=\frac{\epsilon_{f}}{2}\left[(1+\epsilon_{CID})N_{\nu_{f}}+(1-\epsilon_{CID})N_{\bar{\nu}_{f}}\right]$ (1)
$N_{\bar{\nu}_{\text{f,obs}}}$ $=\frac{\epsilon_{f}}{2}\left[(1+\epsilon_{CID})N_{\bar{\nu}_{f}}+(1-\epsilon_{CID})N_{\nu_{f}}\right]$ (2)
with the detection efficiency $\epsilon_{f}$ and $N_{\nu}$ is the number of (anti-)neutrinos of flavor $f$ we get after taking oscillations of the initial flux into account. DUNE could potentially differentiate a muon final state from an anti-muon final state with $\epsilon_{CID}\approx 72\%$ Ternes et al. (2019) from muon capture on argon. In addition, inelasticity can provide some CID information as well.
Larger values of $\epsilon_{CID}$ could be achieved with magnetized detectors, an idea pursued by the INO collaboration. They estimate a CID efficiency of $\epsilon_{CID}\gtrsim 95\%$ for muons with GeV energy Ahmed et al. (2017). Electron CID at GeV energies has been studied in the scenario of a totally active scintillator detector in a magnetic field Bross et al. (2008); Asfandiyarov et al. (2014), and a magnetized LAr detector Rubbia (1977, 2001, 2004, 2009). Given that the final DUNE detector modules remain undesigned, a discussion of the possible benefit of working to improve CID is quite timely. In Huber and Schwetz (2008) it has been pointed out that CID can also be achieved statistically at some level, provided that the energy resolution is sufficiently good.
Our results show that CID can improve the precision on $\delta$ by 15-20%. We have also studied the effect of assuming muon and electron CID at the same time and the results are similar. We find that electron CID has a bigger impact on the improvements in precision on $\delta$. This is likely due to the fact that the detected $\nu_{\mu}\to\nu_{e}$ spectrum is more similar to the detected $\bar{\nu}_{e}\to\bar{\nu}_{e}$ spectrum than the detected $\nu_{e}\to\nu_{\mu}$ spectrum to the detected $\bar{\nu}_{\mu}\to\bar{\nu}_{\mu}$ spectrum, see appendix A. We find that the sensitivity of a neutrino factory comes somewhat more from the $\nu_{e}\to\nu_{\mu}$ golden channel De Rujula et al. (1999); Cervera et al. (2000) and less from the $\nu_{\mu}\to\nu_{e}$ channel which will be leveraged by future proton accelerator neutrino experiments. The oscillation probabilities in these two channels are the same up to the change $\delta\to-\delta$, so each channel serves as an important cross check of the other, but provides no fundamentally new insights into the nature of CP violation or conservation. They do, however, provide a probe of CPT invariance because a neutrino factory is sensitive to four different appearance channels which CP, T, and CPT conjugates of each other. Thus a neutrino factory, especially when combined with DUNE and HK, will provide a valuable unique test of CPT invariance.
Overall, the improvements on precision on $\delta$ with the BNL-SURF setup are larger than at the FNAL-SURF setup largely because in the former setup the neutrinos experience more matter effects Wolfenstein (1978a) due to the higher neutrino energy which increases the appearance probability.
Figure 2: Precision on $\delta$ in degrees as a function of neutrino factory exposure and CID. The neutrino factory has a baseline and muon energy of 1284.9 km and 5 GeV and also includes 10 years of both DUNE and HK. The shaded regions are for electron CID and the lines are for muon CID; electron CID does slightly better in most cases. The horizontal lines show possibly achievable muon CID. The true values of $\delta$ are 00 (left) and $-90^{\circ}$ (right). We assume as ratio of neutrino to anti-neutrino running time of 1:1 for the neutrino factory and DUNE 1:1 and 1:3 for HK.
While increased precision on $\delta$ and other oscillation parameters for the sake of precision is an important goal of particle physics, we also provide some context for these sensitivities. Because $\delta$ is the last remaining oscillation parameter that is largely unconstrained, it plays a key role in flavor model predictions Girardi et al. (2015); Everett et al. (2021), along with other parameters such as $\theta_{12}$ and $\Delta m^{2}_{21}$ Gehrlein et al. (2022) which are expected to become measured with much more precision in the coming years by JUNO Abusleme et al. (2022). While there are many conceivable predictions on neutrino parameters from many different structures (see Denton and Gehrlein (2024) for a recent survey), here we focus on charged lepton corrections as they predict the mixing parameters, but not the masses. In fig. 3 we show the predictions of $\cos\delta$ from four popular charged lepton correction structures as discussed in Everett et al. (2021) alongside the expected $1\sigma$ precision on $\cos\delta$ (taking the wider allowed ranges in all cases due to the sign degeneracy since long-baseline experiments predominantly measure $\sin\delta$ from their appearance channels, although they also probe $\cos\delta$ from their disappearance measurements Denton (2023)). We consider both the precision from the combination of DUNE and HK as well as the expected improvement from including the neutrino factory at the FNAL baseline. In table 1 we document the expected precision on $\delta$ for various experiments.
Figure 3: The predicted regions of $\cos\delta$ due to several popular structures within the charged-lepton correction framework, adapted from Everett et al. (2021); GR1 and GR2 are two flavor model predictions based on the golden ratio Datta et al. (2003); Everett and Stuart (2009); Rodejohann (2009); Adulpravitchai et al. (2009); Feruglio and Paris (2011); Ding et al. (2012), TBM is the tribimaximal flavor model prediction Wolfenstein (1978b); Harrison et al. (2002); Harrison and Scott (2002); Xing et al. (2006); He and Zee (2003), and HEX is based on a hexagonal structure Albright et al. (2010). Along the top are the expected $1\sigma$ sensitivities to $\cos\delta$ at the various central values predicted by the models. DUNE and HK are both at 5+5 years as described in the text, as is neutrino factory with the FNAL baseline.
We see that while after DUNE and HK there will be some modest discriminating capabilities in some cases, the addition of neutrino factory will considerably increase the capability to differentiate flavor models. Combined with JUNO’s constraint on solar parameters should lead to further discrimination capabilities as well.
IV Discussion and Recommendations
Our results show that CID plays a smaller role than emphasized in studies a decade ago. This is primarily due to the fact that LArTPC detectors have sufficiently good energy resolution to distinguish signal from background as well as identify the different flavors. In fact, the LArTPC energy resolution could be even better than assumed here, see e.g. De Romeri et al. (2016); Friedland and Li (2019).
[…]
Using a similar detector setup for a neutrino factory as for near-future experiments could further reduce the uncertainties on the cross section and detector efficiencies, as there is an ongoing program to measure the neutrino cross section on Ar and advancing the LAr technology Machado et al. (2019).
A neutrino factory could be an appealing possible future neutrino oscillation experiment should the results of HK and DUNE disagree. If the tension between NOvA and T2K in the CP violation measurements Denton et al. (2021); Chatterjee and Palazzo (2021) persists, DUNE will be able to probe it Denton et al. (2023). However if the tension requires further oscillation studies at a higher neutrino energy, longer baseline, and overall smaller flux uncertainties than fixed target oscillation experiments a neutrino factory is a favorable setup for future studies. [^1] Note that atmospheric neutrino experiments can also provide oscillation studies at higher energies and longer baselines however their flux uncertainty and flavor composition is more uncertain than at a neutrino factory. Furthermore, at a neutrino factory the neutrino energy is flexible and tunable Delahaye et al. (2013). Finally, at a neutrino factory six oscillation channels and their CP conjugate ones are accessible with similar large number of events as the initial neutrino beam consists equally of $\nu_{e},\nu_{\mu}$.
Table 1: Resolution on $\delta$ for combinations of different experiments. For DUNE we assume 480 kT-MW-yr, for HK we assume 2.47 MT-MW-yr, for the neutrino factory exposure we assume 10 years for a 40 kT LAr detector and $10^{21}$ muon decays per year.
$\delta=(-90^{\circ},~{}0)$ no CID 100% eCID 100% $\mu$CID
HK $(20.8^{\circ},5.6^{\circ})$
DUNE $(17.7^{\circ},9.3^{\circ})$
DUNE+HK $(13.9^{\circ},4.8^{\circ})$
DUNE+HK+NF(FNAL) $(10.6^{\circ},3.8^{\circ})$ $(8.3^{\circ},3.1^{\circ})$ $(8.9^{\circ},3.3^{\circ})$
DUNE+HK+NF(BNL) $(9.2^{\circ},3.9^{\circ})$ $(7.9^{\circ},3.3^{\circ})$ $(8.5^{\circ},3.4^{\circ})$
V Conclusions
Motivated by the recent P5 report, we study the physics potential of a neutrino factory in improving the precision on the determination of the remaining undetermined neutrino oscillation parameters with a focus on the complex phase $\delta$ which governs the amount of CP violation in the neutrino sector. We consider two different baselines and muon energies with a LAr detector based at SURF. Similar baselines and energies were considered as a low-energy neutrino factory Geer et al. (2007); Bross et al. (2008) appropriate to measure CPV for large values of $\theta_{13}$ Tang and Winter (2010); Dighe et al. (2012); Ballett and Pascoli (2012); Christensen et al. (2013). Such a neutrino factory setup could be achieved with a proton beam with 2-4 MW Hol (2000); Oza (2001); Albright et al. (2004); Geer and Zisman (2007); Ankenbrandt et al. (2009).
We find that a neutrino factory has several important benefits for the community:
  • Improved precision on several fundamental parameters including the amount of CP violation,
  • Improved flavor model differentiation capabilities,
  • A technological stepping stone on the way to a high energy muon collider to measure the electroweak sector with increased precision.
Such a machine may also be necessary in the event of a tension in the future oscillation data, a scenario that currently exists with, albeit at low significance. As there are numerous valuable benefits to such a machine, we feel that it would be a strong addition to the neutrino oscillation experimental landscape moving forward.
We also investigated the role of charge identification of electrons or muons (or both). We found that CID can lead to some improvement in the sensitivity to CP violation. The impact is largest with larger statistics as CID helps to cut through systematics which become relevant at high statistics.
A neutrino factory is a natural successor to the upcoming accelerator based long-baseline neutrino oscillation program to begin to move the neutrino sector towards the levels of precision already achieved in other sectors of particle physics.
Acknowledgements.
PBD acknowledges support by the United States Department of Energy under Grant Contract No. DE-SC0012704.
Appendix A Number of events
Figure 4: Number of events at a neutrino factory and DUNE for the BNL-SURF and FNAL-SURF baseline with muon energy 8 GeV and 5 GeV for 5+5 years of running. We show the number of events from the $\nu_{e}\to\nu_{\mu}$ channel in red for varying values of $\delta$ and the dominant background from $\bar{\nu}_{\mu}\to\bar{\nu}_{\mu}$.
In fig. 4 we show the number of events at a neutrino factory and DUNE. The background for the FNAL baseline is larger whereas the signal at the BNL baseline is where the background has a dip. We also note that the shape of the disappearance spectra is quite different than that from a proton fixed target source due to the very different spectral shape, see fig. 1.
Appendix B Results for BNL baseline
In fig. 5 we show the resolution of $\delta=0,-90^{\circ}$ for the baseline between BNL-SURF. We see that a BNL based neutrino factory would lead to significant improvement in the precision on $\delta$, especially for $\delta=-90^{\circ}$, with only a modest benefit from CID.
Figure 5: Resolution on $\delta$ as a function of exposure of a neutrino factory with baseline 2542.3 km, corresponding to the distance between BNL-SURF combined with DUNE and HK. We also show the results for 100% eCID and 100% $\mu$CID as well as without CID.